# The Nuclear Star Cluster of M 74: a fossil record of the very early stages of a star-forming galaxy

### v2

@@
Nuclear star clusters (NSC) are dense and compact stellar systems, of sizes of few parsecs, located at galactic centers. Their properties and formation mechanisms seem to be tightly linked to the evolution of the host galaxy, with potentially different formation channels for late- and early-type galaxies (respectively, LTGs and ETGs). While most observations target ETGs, here we focus on the NSC in M 74 (NGC 628), a relatively massive, gas-rich and star-forming spiral galaxy, part of the PHANGS survey. We analyzed the central arcmin of the PHANGS-MUSE mosaic, in which the NSC is not spatially resolved. We analyzed the NSC stellar populations in a point spread function (PSF) aperture, and compared it to the host galaxy. Within the PSF size, the NSC is contaminated by the host-galaxy light. We performed a two-dimensional spectro-photometric decomposition of the MUSE cube, employing a modified version of the C2D code, to disentangle the NSC from its host. This method provided different data cubes for the NSC and the host galaxy, allowing for both their comparison in a PSF aperture, and the spatially resolved analysis of the host. Our results show a very old and metal-poor NSC, in contrast to the surrounding regions. While similar properties were found in NSCs hosted by galaxies of different masses and/or morphological types from M 74, they are somewhat unexpected for a relatively massive star-forming spiral galaxy. The spatially resolved stellar populations of the host galaxy display much younger (light-weighted) ages and higher metallicities, especially in the central region (${\sim}500$ pc) surrounding the NSC. This suggests that this NSC formed a long time ago, and evolved passively until today, without any further growth. Our results show that the NSC was not involved in the active recent star-formation history of its host galaxy.
@@ 1 Introduction
In the central regions of galaxies, massive black holes often coexist with dense stellar systems known as nuclear star clusters (NSCs). NSCs have sizes of a few parsecs and masses between $10^{4}$ and a few times $10^{8}$ M(e.g., Côté et al. 2006; Georgiev et al. 2016; Sánchez-Janssen et al. 2019; Hoyer et al. 2023a). Their properties trace the dynamical state and evolution of the central regions of galaxies and are key to revealing not only their formation and growth but also the mechanisms funneling stars and gas to the center, leading to the seeding and growth of central black holes (see Neumayer et al. 2020 for a review). NSCs can form via two main channels: star-cluster inspiral to the center due to dynamical friction (e.g., Tremaine et al. 1975; Capuzzo-Dolcetta 1993; Oh and Lin 2000; Lotz et al. 2001; Arca-Sedda and Capuzzo-Dolcetta 2014) and in-situ star formation after gas inflow (e.g., Silk et al. 1987; Mihos and Hernquist 1994; Milosavljević 2004; Bekki et al. 2006; Bekki 2007). These two formation channels are not mutually exclusive, as migrating star clusters can be gas-rich (Guillard et al. 2016), and combinations of in-situ formation and cluster merging have often been necessary to explain the observed properties (Fahrion et al. 2019, 2021, 2022b, 2022a, 2024).
@@
More than 60% of galaxies with stellar masses between $10^{8}$ and $10^{10}$ M host an NSC, with a peak nucleation fraction (${\sim}90$%) near $10^{9}$ M (Sánchez-Janssen et al. 2019; Hoyer et al. 2021). This fraction drops drastically for lower masses. For higher masses, a similar drop is observed for early-type galaxies (ETGs) hosting NSCs, while massive late-type galaxies (LTGs) seem to maintain a similar high fraction (Neumayer et al. 2020; Ashok et al. 2023). This trend, together with NSC properties such as metallicity, suggests a transition in the dominant NSC formation channel at about $10^{9}$ M(e.g. Lyu et al. 2025): many low-mass galaxies host metal-poor NSCs, which have similar properties to globular clusters and are traditionally associated with cluster inspiral (e.g., Fahrion et al. 2020). NSCs are systematically more metal rich at higher galaxy masses, showing significant contributions from younger stellar populations, often ongoing or recent star formation (see, e.g., Fig. 9 in the review from Neumayer et al. 2020, and Fahrion et al. 2021, 2022b, 2022a). However, metal-rich NSCs may also form from the inspiral of young massive star clusters, complicating more the interpretation of observations (Paudel and Yoon 2020; Fahrion et al. 2024).
@@
Most evidence of the trends described above comes from NSCs hosted by ETGs, while it needs to be further investigated in LTGs. Several kinematic and stellar-population studies revealed morphological dependencies. Pinna et al. (2021) analyzed the spatially resolved kinematics of eleven NSCs. While rotation is ubiquitous in the full sample, NSCs in LTGs are typically rotation-dominated and associated with central gas inflow, whereas those in ETGs show slower rotation, more complexity and signatures of past mergers (especially at lower NSC masses). NSCs in very LTGs (Scd or later types) are on average young, but made up of stellar populations of different ages, with prolonged and recurrent star formation until the present day (Walcher et al. 2005, 2006; Rossa et al. 2006; Kacharov et al. 2018). NSCs in LTGs also showed a diversity in their ages and metallicities (Kacharov et al. 2018). Despite the scatter, NSCs follow an average trend being in general older in earlier types spirals (${\sim}1-2$ Gyr) and younger, less massive, and more metal poor in the latest types (about a couple of hundred Myr, Rossa et al. 2006; Walcher et al. 2006). Combined trends with galaxy mass and morphological types support a link between NSC formation and the host-galaxy evolution.
@@
In this work, we present a new version of the spectro-photometric decomposition method C2D that enables the extraction of NSC properties while mapping its surroundings. C2D (Méndez-Abreu et al. 2019) separates an IFS data cube into multiple morphological components, delivering a data cube for each one of them. It uses a very similar approach to BUDDI (see above), with the difference that the latter delivers one spectrum for each morphological component, instead of a full separated data cube as C2D does. This allows us to spatially resolve stellar populations in the NSC surroundings, helping understand the NSC evolution in the context of its host galaxy. We applied this method to the central region of the massive spiral galaxy M 74 (NGC 628) and analyzed its stellar populations. The paper is structured as follows. In Sect. 2, we describe the data set. Section 3 summarizes the known properties of the galaxy M 74 and its NSC. In Sect. 4, we describe the methods used for the spectro-photometric decomposition of the NSC from the host galaxy, and for the stellar-population analysis. The analysis of the non-decomposed IFS data cube is presented in Appendix A. Section 5 describes our results, which are discussed in Sect. 6. Our conclusions are outlined in Sect. 7.
@@ 4 Methods
In this paper, we analyze the stellar populations of the NSC in M 74 in the context of the host galaxy. We first disentangled the NSC light from its host using a spectro-photometric decomposition method, yielding separate data cubes for each component (Sect. 4.1). For the stellar population analysis, we used two overall approaches. One consisted of fitting spectra integrated over a PSF aperture (Sect. 4.1.2), separately for the NSC and the host galaxy (one each, Fig. 4). Secondly, we fitted the spectra of each spaxel in host-galaxy data cube, allowing the spatially resolved analysis of its stellar populations. All the details about the spectral fitting process are given in Sect. 4.2.
@@ 4.1 Spectrophotometric decomposition of the NSC: C2D
As mentioned in Sect. 1, it is important to disentangle the NSC light from the other galaxy components, overlapped in the central region along the LOS, to analyze the properties of the NSC more accurately. For this purpose, we used C2D, a multicomponent spectrophotometric decomposition code (Méndez-Abreu et al. 2019). C2D creates one quasi-monochromatic image for each spectral pixel (i.e. per wavelength), and then performs a multicomponent photometric decomposition of those images. The photometric decomposition of the galaxy surface brightness distribution in each image is done in two dimensions with the code GAlaxy Surface Photometry 2 Dimensional Decomposition (GASP2D, Méndez-Abreu et al. 2008). GASP2D adopts a Levenberg-Marquardt algorithm to estimate the structural parameters that fit the two-dimensional surface brightness of the galaxy. In the first version, it was optimized to fit a bulge and a disk (Méndez-Abreu et al. 2008), while it was later adapted to fit also bars (Méndez-Abreu et al. 2014). In this work, we adapted GASP2D to fit an unresolved NSC, in addition to a bulge and a disk (Sect. 4.1.2).
@@ 4.1.1 Preparation of the data cube of the central region of the galaxy
    Cropping the mosaic. Since this work focuses on the central region of the galaxy, we first cropped the original PHANGS-MUSE mosaic to the central arcmin. This ensured a successful NSC decomposition, avoiding the need to handle the full mosaic, which is computationally expensive. This also avoided offsets in the stellar population parameters between different MUSE pointings, which were detected in PHANGS-MUSE large mosaics due to biases in the sky subtraction (Emsellem et al. 2022). An image obtained from the cropped data cube of the central arcmin, using the full MUSE wavelength range, is shown in Fig. 1. We also created a mask to exclude bright foreground stars from the analysis described below (see, e.g., Fig. 2 and 3).
@@ 4.1.2 Application of C2D to M 74
For this work, we adapted C2D and GASP2D to include the following three components: a NSC, a bulge and a disk. This choice was, on the one hand, motivated by the fact that this galaxy was fitted with a bulge and a disk by Salo et al. (2015) as part of the S4G sample (Sheth et al. 2010). On the other hand, we are mostly interested in the NSC, which is not spatially resolved in MUSE data as its size is smaller than the PSF size (Sect. 3.1). Its light is spread by the PSF in a larger region than its size, and we follow other studies modeling the NSC as a PSF to disentangle its spectra from the galaxy spectra (e.g., Johnston et al. 2020; Fahrion et al. 2021). We fitted the NSC using the Moffat function from the PSF modeling in Emsellem et al. (2022, see Sect. 2 for more details). While we adopted a FWHM of 0.6 arcsec (three MUSE spaxels), we tested that the results were not affected by variations of $\pm 0.2$ arcsec (one spaxel). To test this, we repeated the full decomposition and spectral-fitting process with a FWHM of 0.4 and 0.8 arcsec, finding that the differences in the NSC stellar populations were negligible. The FWHM and the power index (2.8, see Sect. 2) of the Moffat function were fixed in the fitting process, while the PSF (NSC) central intensity was kept free and fitted.
For the bulge and the disk we used respectively a Sérsic and a broken-exponential two-dimensional profiles (see Méndez-Abreu et al. 2017 for the specific implementation). We extensively tested different combinations of free and fixed parameters of these profiles, resulting in a better decomposition with a certain number of them fixed, as described below. Furthermore, the impact of small errors in the bulge and disk parameters on the NSC spectra, which is the goal of the decomposition, is negligible. The effective intensity, the effective radius and the Sérsic parameter of the bulge, as well as the central intensity of the disk, were kept free and fitted by the code, while the ellipticity and the position angle of both the bulge and the disk were fixed to the values from the S4G decomposition by Salo et al. (2015), who fitted the disk with a single exponential. We checked that the effective radius and the Sérsic parameter of the bulge resulting from the fit were very similar to the values from Salo et al. (2015). We did not expect to find the same values since they used a different wavelength range compared to ours.
@@
Since we are interested in the stellar-population properties of the NSC and its comparison with the surrounding host galaxy, we will use for the rest of the analysis a data cube for the NSC and another for the rest of the galaxy. The latter was obtained by combining the bulge and the disk data cubes, and corresponds to the host galaxy, i.e. the galaxy after subtracting the NSC. We preferred this approach instead of considering separately the bulge and the disk, to minimize the impact of a potential suboptimal decomposition of these two morphological structures. This also follows previous studies of this kind, comparing the NSC to the host galaxy as a whole (Johnston et al. 2020; Fahrion et al. 2021). However, the C2D approach, providing spatially resolved spectral information, additionally allows us to fairly compare NSC properties with the host galaxy in the same aperture. This minimizes biases derived from quantitative comparisons of different spatial regions, due to spatial gradients in the host galaxy.
@@
Table 1: Average stellar-population properties and their uncertainties, obtained as detailed in Sect. 4.2 from spectra integrated in a $1\sigma$ PSF aperture. From top to bottom: the NSC, the host galaxy (decomposed as explained in Sect. 4.1), and all components combined in the LOS (obtained from the original data cube). From left to right, columns indicate: light-weighted (LW) average age, total metallicity [M/H] and [Mg/Fe] abundance; mass-weighted (MW) average age, [M/H] and [Mg/Fe]. Uncertainties were obtained via Monte Carlo simulations (Sect. 4.2).
@@ 4.2 Spectral fitting
We used in this work the semi-empirical sMILES SSP models (Knowles et al. 2023), the most up-to-date version of the MILES SSP models (Vazdekis et al. 2010, 2015). sMILES models are based on BaSTI isochrones (Pietrinferni et al. 2004, 2006) and offer five different values of the [Mg/Fe] abundances, covering the range between -0.2 and 0.6 dex (in steps of 0.2 dex). Throughout this paper, we will use [Mg/Fe] as a proxy for the enhancement in $\alpha$ elements. sMILES models cover ages between 0.03 and 14 Gyr, with 53 different values, and ten values of total metallicities [M/H] between -1.79 and 0.26 dex. This gives a total of 2650 SSP models. Since each of them corresponds to one combination of age, [M/H] and [Mg/Fe], coupling these models with pPXF allows us to extract the distribution of these three stellar-population parameters from each spectrum (given by the distribution of weights given to the different SSPs). These models cover a wavelength range between 3540 and 7409.6 $\mathrm{\SIUnitSymbolAngstrom}$, at a spectral resolution of 2.51 $\mathrm{\SIUnitSymbolAngstrom}$ at FWHM. Among sMILES sets of models, we chose the set of SSPs that adopted a bimodal initial mass function with a logarithmic slope of 1.3 for the segment of massive stars (Vazdekis et al. 2003). These models were normalized to correspond to one solar mass each, so that mass-weighted results are obtained from pPXF if SSPs are not renormalized. They need to be renormalized to their own median to obtain light-weighted results. We used in this work both normalizations, to obtain light-weighted stellar population maps and average results, and mass-weighted results to allow us to reconstruct SFHs in terms of mass fractions corresponding to different stellar populations.
@@
Our general approach for the stellar-population analysis was to average results from Monte Carlo simulations of unregularized pPXF fits, following Cappellari (2023). We used this approach to fit the three spectra extracted from the PSF aperture (Sect. 4.1.2), i.e. the NSC, the host-galaxy, and for comparison the spectrum from the original integrated data cube. We performed 100 Monte Carlo realizations for each spectrum. For each realization, we perturbed the spectrum with random, wavelength-dependent noise drawn from a Gaussian distribution. The noise level was taken from the residuals at the specific wavelength, from an initial (unregularized) pPXF fit. From the distribution of SSP weights for the 100 realizations, we estimated the mean ages, [M/H] and [Mg/Fe], and their statistical uncertainties (as standard deviations), for each of the three original spectra. This allowed us not to use any regularization on the SSP weights during the pPXF fits.
@@
The resulting host-galaxy data cube from the decomposition was rebinned to a higher S/N of 100 before performing the spectral fitting for the spatially resolved stellar population analysis, ensuring robust results and lower uncertainties in the outer fainter region. This still left us with 6766 Voronoi bins, which made the Monte Carlo approach time-consuming and computationally expensive for the entire data cube. Thus, we decided to map the stellar-population parameters of the host galaxy using pPXF regularized fits with a unit regularization parameter. While one single unregularized fit gives discrete and noisy solutions, some level of regularization ensures smoother and more realistic stellar-population solutions (e.g., Cappellari 2017). However, regularization above a certain level can lead to biases in the results and dilute features in the SFHs. Thus, we used the approach presented in Pinna et al. (2019b) to choose the regularization parameter (described as follows, see also Pinna et al. 2019a; Martig et al. 2021; Sattler et al. 2023, 2025). After rescaling the pPXF noise parameter as the one necessary to obtain a reduced $\chi^{2}=1$ with no regularization, we choose a regularization well below a certain maximum calibration value. We defined the calibration regularization as the minimum value increasing $\chi^{2}$, from the case with no regularization, by $\Delta\chi^{2}\approx\sqrt{2N_{\text{goodpix}}}$, where $N_{\text{goodpix}}$ is the number of fitted (spectral) pixels (McDermid et al. 2015; Cappellari 2017; Boecker et al. 2020). For comparison, we also mapped the stellar populations from the original data cube before the C2D decomposition, using this same approach (S/N=100=100, multiplicative polynomial of degree 8 and unit regularization parameter).

### v3

# arXiv:2512.03999v3[astro-ph.GA] 02 Mar 2026

@@
Nuclear star clusters (NSC) are dense and compact stellar systems, of sizes of few parsecs, located at galactic centers. Their properties and formation mechanisms seem to be tightly linked to the evolution of the host galaxy, with potentially different formation channels for late- and early-type galaxies (respectively, LTGs and ETGs). While most observations target ETGs, here we focus on the NSC in M 74 (NGC 628), a relatively massive, gas-rich and star-forming spiral galaxy, part of the PHANGS survey. We analyzed the central arcmin of the PHANGS-MUSE mosaic, in which the NSC is not spatially resolved. We analyzed the NSC stellar populations in a point spread function (PSF) aperture, and compared it to the host galaxy. Within the PSF size, the NSC is contaminated by the host-galaxy light. We performed a two-dimensional spectro-photometric decomposition of the MUSE cube, employing a modified version of the C2D code, to disentangle the NSC from its host. This method provided different data cubes for the NSC and the host galaxy, allowing for both their comparison in a PSF aperture, and the spatially resolved analysis of the host. Our results show a very old and metal-poor NSC, in contrast to the surrounding regions. While similar properties were found in NSCs hosted by galaxies of different masses and/or morphological types from M 74, they are somewhat unexpected for a relatively massive star-forming spiral galaxy. The spatially resolved stellar populations of the host galaxy display much younger (light-weighted) ages and higher metallicities, especially in the central region (${\sim}500$ pc) surrounding the NSC. This suggests that this NSC formed a long time ago, and evolved passively until today, without any further growth. No significant amounts of gas would have reached the very central region in the last 8 Gyr.
@@ 1 Introduction
In the central regions of galaxies, massive black holes often coexist with dense stellar systems known as nuclear star clusters (NSCs). NSCs have sizes of a few parsecs and masses between $10^{4}$ and a few times $10^{8}$ M(e.g., Böker et al. 2002, 2004; Walcher et al. 2005; Côté et al. 2006; Georgiev et al. 2016; Sánchez-Janssen et al. 2019; Hoyer et al. 2023a). Their properties trace the dynamical state and evolution of the central regions of galaxies and are key to revealing not only their formation and growth but also the mechanisms funneling stars and gas to the center, leading to the seeding and growth of central black holes (see Neumayer et al. 2020 for a review). NSCs can form via two main channels: star-cluster inspiral to the center due to dynamical friction (e.g., Tremaine et al. 1975; Capuzzo-Dolcetta 1993; Oh and Lin 2000; Lotz et al. 2001; Arca-Sedda and Capuzzo-Dolcetta 2014) and in-situ star formation after gas inflow (e.g., Silk et al. 1987; Mihos and Hernquist 1994; Milosavljević 2004; Bekki et al. 2006; Bekki 2007). These two formation channels are not mutually exclusive, as migrating star clusters can be gas-rich (Guillard et al. 2016), and combinations of in-situ formation and cluster merging have often been necessary to explain the observed properties (Fahrion et al. 2019, 2021, 2022b, 2022a, 2024).
@@
More than 60% of galaxies with stellar masses between $10^{8}$ and $10^{10}$ M host an NSC, with a peak nucleation fraction (${\sim}90$%) near $10^{9}$ M (Sánchez-Janssen et al. 2019; Hoyer et al. 2021). This fraction drops drastically for lower masses. For higher masses, a similar drop is observed for early-type galaxies (ETGs) hosting NSCs, while massive late-type galaxies (LTGs) seem to maintain a similar high fraction (Neumayer et al. 2020; Ashok et al. 2023). This trend, together with NSC properties such as metallicity, suggests a transition in the dominant NSC formation channel at about $10^{9}$ M(e.g. Lyu et al. 2025): many low-mass galaxies host metal-poor NSCs, which have similar properties to globular clusters and are traditionally associated with cluster inspiral (e.g., Fahrion et al. 2020). NSCs are systematically more metal-rich at higher galaxy masses, showing significant contributions from younger stellar populations, often ongoing or recent star formation (see, e.g., Fig. 9 in the review from Neumayer et al. 2020, and Fahrion et al. 2021, 2022b, 2022a). However, metal-rich NSCs may also form from the inspiral of young massive star clusters, complicating the interpretation of observations (Paudel and Yoon 2020; Fahrion et al. 2024).
@@
Most evidence of the trends described above comes from NSCs hosted by ETGs, while it needs to be further investigated in LTGs. Several kinematic and stellar-population studies revealed morphological dependencies. Pinna et al. (2021) analyzed the spatially resolved kinematics of eleven NSCs. While rotation is ubiquitous in the full sample, NSCs in LTGs are typically rotation-dominated and associated with central gas inflow, whereas those in ETGs show slower rotation, more complexity and signatures of past mergers (especially at lower NSC masses). NSCs in very LTGs (Scd or later types) are on average young, but made up of stellar populations of different ages, with prolonged and recurrent star formation until the present day (Walcher et al. 2005, 2006; Rossa et al. 2006; Kacharov et al. 2018). NSCs in LTGs also showed a diversity in their ages and metallicities (Kacharov et al. 2018). Despite the scatter, NSCs follow an average trend being generally older in earlier-type spirals (${\sim}1-2$ Gyr) and younger, less massive, and more metal-poor in the latest types (about a couple of hundred Myr, Rossa et al. 2006; Walcher et al. 2006). Combined trends with galaxy mass and morphological types support a link between NSC formation and the host-galaxy evolution.
@@
In this work, we present a new version of the spectro-photometric decomposition method C2D that enables the extraction of the stellar populations of the NSC while mapping its surroundings. C2D (Méndez-Abreu et al. 2019) separates an IFS data cube into multiple morphological components, delivering a data cube for each one of them. It uses a very similar approach to BUDDI (see above), with the difference that the latter delivers one spectrum for each morphological component, instead of a full, separated data cube as C2D does. This allows us to spatially resolve stellar populations in the NSC surroundings, helping understand the NSC evolution in the context of its host galaxy. We applied this method to the central region of the massive spiral galaxy M 74 (NGC 628) and analyzed its stellar populations. The paper is structured as follows. In Sect. 2, we describe the data set. Section 3 summarizes the known properties of the galaxy M 74 and its NSC. In Sect. 4, we describe the methods used for the spectro-photometric decomposition of the NSC from the host galaxy, and for the stellar-population analysis. The analysis of the non-decomposed IFS data cube is presented in Appendix A. Section 5 describes our results, which are discussed in Sect. 6. Our conclusions are outlined in Sect. 7.
@@ 4 Methods
The goal of this paper is to analyze the stellar populations of the NSC in M 74 in the context of the host galaxy, while we are not interested in the kinematic properties (they would be unresolved for the NSC). We first disentangled the NSC light from its host using a spectro-photometric decomposition method, yielding separate data cubes for each component (Sect. 4.1). For the stellar population analysis, we used two overall approaches. One consisted of fitting spectra integrated over a PSF aperture (Sect. 4.1.2), separately for the NSC and the host galaxy (one each, Fig. 4). Secondly, we fitted the spectra of each spaxel in the host-galaxy data cube, allowing the spatially resolved analysis of its stellar populations. All the details about the spectral fitting process are given in Sect. 4.2.
@@ 4.1 Spectrophotometric decomposition of the NSC: C2D
As mentioned in Sect. 1, disentangling the NSC light from the other galaxy components, overlapped in the central region along the LOS, is essential to analyze the properties of the NSC more accurately. For this purpose, we used C2D, a multicomponent spectrophotometric decomposition code (Méndez-Abreu et al. 2019). C2D creates one quasi-monochromatic image for each spectral resolution element, and then performs a multicomponent photometric decomposition of those images. The photometric decomposition of the galaxy surface brightness distribution in each image is done in two dimensions with the code GAlaxy Surface Photometry 2 Dimensional Decomposition (GASP2D, Méndez-Abreu et al. 2008). GASP2D adopts a Levenberg-Marquardt algorithm to estimate the structural parameters that fit the two-dimensional surface brightness of the galaxy. In the first version, it was optimized to fit a bulge and a disk (Méndez-Abreu et al. 2008), while it was later adapted to fit also bars (Méndez-Abreu et al. 2014). In this work, we adapted GASP2D to fit an unresolved NSC, in addition to a bulge and a disk (Sect. 4.1.2).
@@ 4.1.1 Preparation of the data cube of the central region of the galaxy
    Cropping the mosaic. Since this work focuses on the central region of the galaxy, we first cropped the PHANGS-MUSE mosaic to the central arcmin. This ensured a successful NSC decomposition, avoiding the need to handle the full mosaic, which is computationally expensive. This also avoided offsets in the stellar population parameters between different MUSE pointings, which were detected in PHANGS-MUSE large mosaics due to biases in the sky subtraction (Emsellem et al. 2022). An image obtained from the cropped data cube of the central arcmin, using the full MUSE wavelength range, is shown in Fig. 1. We also created a mask to exclude bright foreground stars from the analysis described below (see, e.g., Fig. 2 and 3).
@@ 4.1.2 Application of C2D to M 74
For this work, we adapted C2D and GASP2D to include the following three components: an NSC, a bulge and a disk. This choice was, on the one hand, motivated by the fact that this galaxy was fitted with a bulge and a disk by Salo et al. (2015) as part of the S4G sample (Sheth et al. 2010). On the other hand, we are mostly interested in the NSC, which is not spatially resolved in MUSE data as its size is smaller than the PSF size (Sect. 3.1). Its light is spread by the PSF in a larger region than its size, and we follow other studies modeling the NSC as a PSF to disentangle its spectra from the galaxy spectra (e.g., Johnston et al. 2020; Fahrion et al. 2021). We fitted the NSC using the Moffat function from the PSF modeling in Emsellem et al. (2022, see Sect. 2 for more details). While we adopted a FWHM of 0.6 arcsec (three MUSE spaxels), we tested that the results were not affected by variations of $\pm 0.2$ arcsec (one spaxel). To test this, we repeated the full decomposition and spectral-fitting process with a FWHM of 0.4 and 0.8 arcsec, finding that the differences in the NSC stellar populations were negligible. The FWHM and the power index (2.8, see Sect. 2) of the Moffat function were fixed in the fitting process, while the PSF (NSC) central intensity was kept free and fitted.
For the bulge and the disk, we used a Sérsic and a broken-exponential two-dimensional profiles, respectively (see Méndez-Abreu et al. 2017 for the specific implementation). We extensively tested different combinations of free and fixed parameters of these profiles, resulting in a better decomposition with a certain number of them fixed, as described below. Furthermore, the impact of small errors in the bulge and disk parameters on the NSC spectra, which is the goal of the decomposition, is negligible. The effective intensity, the effective radius and the Sérsic parameter of the bulge, as well as the central intensity of the disk, were kept free and fitted by the code, while the ellipticity and the position angle of both the bulge and the disk were fixed to the values from the S4G decomposition by Salo et al. (2015), who fitted the disk with a single exponential. We checked that the effective radius and the Sérsic parameter of the bulge resulting from the fit were very similar to the values from Salo et al. (2015). We did not expect to find the same values since they used a different wavelength range compared to ours.
@@
Since we are interested in the stellar-population properties of the NSC and its comparison with the surrounding host galaxy, we will use for the rest of the analysis a data cube for the NSC and another for the rest of the galaxy. The latter was obtained by combining the bulge and the disk data cubes, and corresponds to the host galaxy, i.e. the galaxy after subtracting the NSC. We preferred this approach instead of considering the bulge and the disk separately to minimize the impact of a potential suboptimal decomposition of these two morphological structures. This also follows previous studies of this kind, comparing the NSC to the host galaxy as a whole (Johnston et al. 2020; Fahrion et al. 2021). However, the C2D approach, providing spatially resolved spectral information, additionally allows us to fairly compare NSC properties with the host galaxy in the same aperture. This minimizes biases derived from quantitative comparisons of different spatial regions, due to spatial gradients in the host galaxy.
@@
Table 1: Average stellar-population properties and their uncertainties, obtained as detailed in Sect. 4.2 from spectra integrated in a $1\sigma$ PSF aperture. From top to bottom: the NSC, the host galaxy (decomposed as explained in Sect. 4.1), and all components combined in the LOS (obtained from the original - kinematic-corrected - data cube). From left to right, columns indicate: light-weighted (LW) average age, total metallicity [M/H] and [Mg/Fe] abundance; mass-weighted (MW) average age, [M/H] and [Mg/Fe]. Uncertainties were obtained via Monte Carlo simulations (Sect. 4.2).
@@ 4.2 Spectral fitting
We used in this work the semi-empirical sMILES SSP models (Knowles et al. 2023), the most up-to-date version of the MILES SSP models (Vazdekis et al. 2010, 2015). sMILES models are based on BaSTI isochrones (Pietrinferni et al. 2004, 2006) and offer five different values of the [Mg/Fe] abundances, covering the range between -0.2 and 0.6 dex (in steps of 0.2 dex). Throughout this paper, we will use [Mg/Fe] as a proxy for the enhancement in $\alpha$ elements. sMILES models cover ages between 0.03 and 14 Gyr, with 53 different values, and ten values of total metallicities [M/H] between -1.79 and 0.26 dex. This gives a total of 2650 SSP models. Since each of them corresponds to one combination of age, [M/H] and [Mg/Fe], coupling these models with pPXF allows us to extract the distribution of these three stellar-population parameters from each spectrum (given by the distribution of weights given to the different SSPs). These models cover a wavelength range between 3540 and 7409.6 $\mathrm{\SIUnitSymbolAngstrom}$, at a spectral resolution of 2.51 $\mathrm{\SIUnitSymbolAngstrom}$ at FWHM. Among sMILES sets of models, we chose the set of SSPs that adopted a bimodal initial mass function with a logarithmic slope of 1.3 for the segment of massive stars (Vazdekis et al. 2003). These models were normalized to correspond to one solar mass each, so that mass-weighted results are obtained from pPXF if SSPs are not renormalized. They need to be renormalized to their own median to obtain light-weighted results. We used both normalizations in this work, to obtain light-weighted stellar population maps and average results, and mass-weighted results to allow us to reconstruct SFHs in terms of mass fractions corresponding to different stellar populations.
@@
Our general approach for the stellar-population analysis was to average results from Monte Carlo simulations of unregularized pPXF fits, following Cappellari (2023). We used this approach to fit the three spectra extracted from the PSF aperture (Sect. 4.1.2), i.e. the NSC, the host-galaxy, and for comparison the spectrum from the original integrated data cube. We performed 100 Monte Carlo realizations for each spectrum. For each realization, we perturbed the spectrum with random, wavelength-dependent noise drawn from a Gaussian distribution. The noise level was taken from the residuals at the specific wavelength, from an initial (unregularized) pPXF fit. From the distribution of SSP weights for the 100 realizations, we estimated the mean ages, [M/H] and [Mg/Fe], and their statistical uncertainties (as standard deviations), for each of the three spectra. This allowed us not to use any regularization on the SSP weights during the pPXF fits.
@@
The resulting host-galaxy data cube from the decomposition was rebinned to a higher S/N of 100 before performing the spectral fitting for the spatially resolved stellar population analysis, ensuring robust results and lower uncertainties in the outer, fainter region. This still left us with 6766 Voronoi bins, which made the Monte Carlo approach time-consuming and computationally expensive for the entire data cube. Thus, we decided to map the stellar-population parameters of the host galaxy using pPXF regularized fits with a unit regularization parameter. While one single unregularized fit gives discrete and noisy solutions, some level of regularization ensures smoother and more realistic stellar-population solutions (e.g., Cappellari 2017). However, regularization above a certain level can lead to biases in the results and dilute features in the SFHs. Thus, we used the approach presented in Pinna et al. (2019b) to choose the regularization parameter (described as follows, see also Pinna et al. 2019a; Martig et al. 2021; Sattler et al. 2023, 2025). After rescaling the pPXF noise parameter as the one necessary to obtain a reduced $\chi^{2}=1$ with no regularization, we choose a regularization well below a certain maximum calibration value. We defined the calibration regularization as the minimum value increasing $\chi^{2}$, from the case with no regularization, by $\Delta\chi^{2}\approx\sqrt{2N_{\text{goodpix}}}$, where $N_{\text{goodpix}}$ is the number of fitted (spectral) pixels (McDermid et al. 2015; Cappellari 2017; Boecker et al. 2020). This approach was also used to map the stellar populations from the original data cube before the C2D decomposition (Appendix A).